OpenStudy (hjkhjkhjkhjk):

The number of deer in a wildlife preserve is given by the function D(t)=184(1.1)0.5tD(t)=184(1.1)0.5t, where tt is time in years. How long does it take the deer population to triple? (You may use a calculator for this problem)

OpenStudy (hjkhjkhjkhjk):

Select one: a. 20.8 years b. 23.1 years c. 24.6 years d. 32.1 years

OpenStudy (joannablackwelder):

Do you have any thoughts on how to start?

OpenStudy (hjkhjkhjkhjk):

Sorry, it came out wrong: D(t)=184(1.1)^0.5t

OpenStudy (hjkhjkhjkhjk):

I think it's C.

OpenStudy (hjkhjkhjkhjk):

I tried plugging in, but it doesn't seem perfect.

OpenStudy (joannablackwelder):

How did you plug it in?

OpenStudy (joannablackwelder):

I agree that C is close, but I think there is a better answer.

OpenStudy (joannablackwelder):

Any other thoughts?

OpenStudy (mathmale):

For starters, please indicate exponentiation correctly: Correct: D(t)=184(1.1)^(0.5t) Incorrect: D(t)=184(1.1)0.5t

OpenStudy (mathmale):

"Triple" means that the number of deer is no longer D(t), but 3[D(t)]. Thus, you need to solve 3D(t) =3 [ 184(1.1)^(0.5t) ] for t.